one number is 3 more than another. the smaller number decreased by 4 times the larger number is the same as 12 less than six times the larger number. find teh larger number.
Answer by edjones(8007)

x=y+3
y-4x=6x-12
y-4(y+3)=6(y+3)-12
y-4y-12=6y+18-12
-3y-12=6y+6
-9y=18
y=-2, x=1
.
Ed